OpenStudy (anonymous):

Lisa invests $4,000 in two types of bonds: X, which offers a 10% return, and Y, which offers a 6% return. If the combined return on the investment is $300, how much did Lisa invest at the rate of 6%? $3,000 $2,500 $1,500 $500

OpenStudy (anonymous):

im so lost...... sorry

OpenStudy (jack1):

will this work? so like 4000 = X+Y X * 0.1 + Y * 0.06 = 300

OpenStudy (anonymous):

idk what your saying thtas why im lost idk what the question wants me to do ;-;

OpenStudy (anonymous):

i dont know how to get the anwser

OpenStudy (jack1):

yeah, that makes sense, then the answer you want is Y... yeah? OK, so X = amount money Lisa invests at 10% Y = amount of money Lisa invests at 6% so total amount invested is $4000... with me so far?

OpenStudy (anonymous):

i guess yeah im with you for now ...lol

OpenStudy (jack1):

cool cool, so next bit the amount of profit she made was $300 so $300 = X times 10% + Y times 6% $300 = X times 0.10 + Y times 0.06 still with me?

OpenStudy (anonymous):

i see what your saying but idk how you got all that like ....im kinda geting it but then im lost @_@

OpenStudy (anonymous):

@-@

OpenStudy (jack1):

welllll, just lets see if this is right firstm then come back and explain if that;s cool?

OpenStudy (anonymous):

yeah but whats the rite anwser??

OpenStudy (jack1):

ok so X + Y = 4000 as she invested $4000 and 0.1X + 0.06Y = 300 as she made $300 profit so rearrange X + Y = 4000 Y = 4000 - X now sub this into the other equation 0.1X + 0.06Y = 300 0.1X + 0.06(4000 - X) = 300 0.1X + 240 - 0.06X) = 300 0.04X = 60 X = 60 / 0.04 = 1500 so if X + Y = 4000 and X = 1500 then Y = 2500

OpenStudy (jack1):

so lisa invested 2500 @ 6% interest rate

OpenStudy (anonymous):

oh ok i see what your doing well klinda but im get stuck on the rearanging parnt i know....but ima try !

OpenStudy (jack1):

sweet, i totally confused everything at the very start too, sorry